# Using convolutional neural networks to detect edge localized modes in DIII-D from Doppler backscattering measurements

### v2

@@ II.1 Input Data
Raw DBS data is a time-series waveform with real and complex components from eight channels, each operating at a different microwave frequency [17]. Channels with higher probe frequencies penetrate further into the plasma, measuring perturbations at different depths within the pedestal. Only three channels were used for training and testing, channels 1 (55 GHz), 4 (62.5 GHz), and 7 (72.5 GHz), to minimize model size. This data was preprocessed before being fed to the model. The preprocessing involves performing a discrete fast Fourier transform to the waveform to obtain a power spectrum density (PSD), where each time step contains data of the signal power as a function of frequency. We then smoothed the data by performing a running mean over the frequency domain of each time step as the model learned faster after smoothing. The preprocessed data is a 3D array with dimensions of time, channel (probe beam frequency), and Doppler shifted frequency.
@@ III.2 Training
To train our model we used DBS and Dα data from three shots — 170869, 170870, and 170878 — which were part of an experimental effort to understand the inter-ELM turbulence behavior [22] and were recorded at a sampling rate of 8 MHz, as opposed to the routine 5 MHz. These discharges have input power very close to the L-H power threshold and hence have nearly evenly spaced large type-I ELMs.

### v3

@@ II.1 Input Data
Raw DBS data is a time-series waveform with real and complex components from eight channels, each operating at a different microwave frequency [17]. Channels with higher probe frequencies penetrate further into the plasma, measuring perturbations at different depths within the pedestal. Only three channels were used for training and testing, channels 1 (55 GHz), 4 (62.5 GHz), and 7 (72.5 GHz), to minimize model size. This data was preprocessed before being fed to the model. The preprocessing involves performing a discrete fast Fourier transform to the waveform to obtain a power spectrum density (PSD), where each time step contains data of the signal power as a function of Doppler shifted frequency. We then smoothed the data by performing a running mean over the frequency domain of each time step as the model learned faster after smoothing. The preprocessed data is a 3D array with dimensions of time, channel (probe beam frequency), and Doppler shifted frequency.
@@ III.2 Training
To train our model, we used DBS and Dα data from three shots — 170869, 170870, and 170878 — which were part of an experimental effort to understand the inter-ELM turbulence behavior [22] and were recorded at a sampling rate of 8 MHz, as opposed to the routine 5 MHz. These discharges have input power very close to the L-H power threshold and hence have nearly evenly spaced large type-I ELMs.